OpenStudy (ellie202000):

I will medal the best response. A satellite can move at 17,000 miles per hour. How many hours will it take to reach a star that is three light years away? Recall that 1 light year = 9.5x10^12 kilometers and 1 kilometer = 0.62 miles. Show your work.

OpenStudy (anonymous):

I have no idea how to solve this, but i have a black/tan doxie!!!!

OpenStudy (ellie202000):

Really?!?! I have a dapple and tan girl, a black and tan girl, and a black and cream boy. And they are all so stubborn.

OpenStudy (anonymous):

I sadly one have the black/tan one, and he's a boy named cooper, he is turning 5 in a few days :)

OpenStudy (ellie202000):

Copper? that's a cute name. Mine girls are turning two in may and my boy is turning one in may. Their names are Piper, Rosie, and Dexter.

OpenStudy (anonymous):

Cool! My mom wants to get a dog named Dexter. You are lucky to have a puppy! :)

OpenStudy (ellie202000):

He is so stubborn though. He like to nip at the girl until they cant stand it anymore and they bite him... then he does it again. He also likes to be held and rests his head on my shoulder. He is just the biggest baby you will ever meet.

OpenStudy (anonymous):

Yeah, I also have a vizsla, which is a bigger dog, and my dachshund likes to boss him around, he thinks he's in charge.

OpenStudy (ellie202000):

We also have a Pitbull/Ikeda who thinks she is a dachshund and tries to lay on our laps.

OpenStudy (anonymous):

do your doxies like to stare at the window for hours? Mine will stand on the window ledge looking at birds, wagging his tail.

OpenStudy (ellie202000):

No and they don't like cars either. Only one like cars and parks but the other two freak out if they know they are going to the park.

OpenStudy (ellie202000):

Here are pictures and a video of my dogs.

OpenStudy (anonymous):

Mine is scared of any other dogs except our other one, Bodi.

OpenStudy (anonymous):

Woah, is the one that's your profile pic blind in one eye?

OpenStudy (ellie202000):

no she just has two different colored eyes.

OpenStudy (ellie202000):

Copper is sooo cute.

OpenStudy (anonymous):

Oh, I when I was a baby, we had a brown doxie named Teapot, and she was blind in one eye.

OpenStudy (ellie202000):

Oh. I think her two different colored eyes does affect her eyes a little bit but not nearly enough to consider her blind.

OpenStudy (anonymous):

Oh, do you have squirrels where you live? Mine is obsessed with chasing squirrels. He has no chance of catching one.

OpenStudy (ellie202000):

Ya but we have more birds like geese and the one with two different colored eyes think she can kill a goose but when she get close she decides the goose can kill her and runs off. They are also obsessed with cats they just want to play with them luckily they don't want to hurt them.

OpenStudy (anonymous):

do you have cats?

OpenStudy (anonymous):

I have a hamster and the doxie wants to kill her.

OpenStudy (ellie202000):

We used to have a hamster and the black and tan one would stare at it for hours.

OpenStudy (anonymous):

We have a blue inflatable ball that looks like that hamster ball, so cooper stalks it, thinking its in it

OpenStudy (anonymous):

we keep the cage on a dresser so the dogs cant get it

OpenStudy (anonymous):

are you there?

OpenStudy (ellie202000):

Ya.

OpenStudy (ellie202000):

We had our hamster on the ground.

OpenStudy (anonymous):

your dogs never got it?

OpenStudy (anonymous):

you should bump your question so someone who knows how to do it can help you, cuz i have no idea

OpenStudy (anonymous):

Is is algebra 2?

OpenStudy (ellie202000):

No they never got it. They're not that smart. And not it is algebra 1.

OpenStudy (anonymous):

I am in algebra 1.

OpenStudy (ellie202000):

Really the lesson is called Scientific Notations.

OpenStudy (ellie202000):

Scientific notations aren't that hard but this problem is confusing.

OpenStudy (anonymous):

i think that maybe if you figure out how many miles per lightyear, then you might be able to figure out the rest easily/

OpenStudy (ellie202000):

Well I know that 17,000miles= 27,358.848 kilometers and 3 light year = 28,500,000,000,000 kilometers but after that I am lost.

OpenStudy (ellie202000):

I would think you would divide them but the answer is 1.0417105281626e+9

OpenStudy (ellie202000):

So would it be one hour?

OpenStudy (ellie202000):

But how can it be one hour if the satellite travels 27,358.848 kilometers in one hour and it has to travel 28,500,000,000,000 kilometers?

OpenStudy (anonymous):

idk, here ill tag some people to help @AlexandervonHumboldt2 @TheEdwardsFamily

OpenStudy (theedwardsfamily):

idk this

OpenStudy (ellie202000):

@ganeshie8

OpenStudy (ellie202000):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

are you able to convert 3 lightyears to miles?

OpenStudy (ellie202000):

well 1 light year is \[5.87849981\times10^{12}\]miles. So the formula for 3 light years is \[3(5.87849981\times10^{12})\] Maybe? I am not sure.

jimthompson5910 (jim_thompson5910):

Let's use what they give us "Recall that 1 light year = 9.5x10^12 kilometers and 1 kilometer = 0.62 miles"

jimthompson5910 (jim_thompson5910):

LY = light year 1 LY * (9.5x10^12 km/1 LY) * (0.62 mi/1 km) = 5.89 x 10^12

jimthompson5910 (jim_thompson5910):

So 1 lightyear = 5.89 x 10^12 miles (approx)

jimthompson5910 (jim_thompson5910):

3 lightyears would be 1 lightyear = 5.89 x 10^12 miles 3*1 lightyear = 3*5.89 x 10^12 miles 3 lightyear = 17.67 x 10^12 miles 3 lightyear = 1.767 x 10^1 x 10^12 miles 3 lightyear = 1.767 x 10^13 miles with me so far?

OpenStudy (ellie202000):

yep

jimthompson5910 (jim_thompson5910):

Speed = 17,000 miles per hour = 1.7 x 10^4 mph

jimthompson5910 (jim_thompson5910):

Distance = Rate * Time (1.767 x 10^13 miles) = (1.7 x 10^4 mph)*(T) solve for T to get T = (1.767 x 10^13)/(1.7 x 10^4) T = ??

OpenStudy (ellie202000):

1.0394117647059e+17?

jimthompson5910 (jim_thompson5910):

incorrect

OpenStudy (phi):

when you divide you subtract the exponents

OpenStudy (ellie202000):

1039411764.71?

jimthompson5910 (jim_thompson5910):

1039411764.71 = 1.03941176471 x 10^9 correct

jimthompson5910 (jim_thompson5910):

1.03941176471 x 10^9 hours is essentially 1 billion hours (a bit more)

OpenStudy (ellie202000):

So the satellite had to travel 1.03941176471x10^9=1039411764.71 hours?

jimthompson5910 (jim_thompson5910):

roughly, yes

OpenStudy (ellie202000):

Ok thank you guys so much.

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

hi